OpenStudy (anonymous):

A missile guidance system has five fail-safe components. The probability of each failing is 0.05. Find these probabilities

OpenStudy (lgbasallote):

do you mean probability that all 5 will fail? then it will be \[0.05 \times 0.05 \times 0.05 \times 0.05 \times 0.05\]

OpenStudy (kropot72):

What are the probabilities required? The question is not complete.

OpenStudy (anonymous):

A missile guidance system has five fail-safe components. The probability of each failing is 0.30. Find the probability that exactly four will fail

OpenStudy (kropot72):

P(4 fail) = 0.3 * 0.3 * 0.3 * 0.3

OpenStudy (lgbasallote):

your question changed.o.O

OpenStudy (zarkon):

5*.3 * .3 * .3 * .3*.7

OpenStudy (zarkon):

assuming independence

OpenStudy (kropot72):

Because each component is fail-safe, the failure of any one of the 5 components does not affect the others. The compound event 4 failures is made up of 4 separate and independent sub-events. Therefore the probability of the compound event is the product of the separate probabilities of each sub-event: P(4 fail) = 0.3 * 0.3 * 0.3 * 0.3

OpenStudy (zarkon):

remember there are 5 components...so to get exactly 4 failures you need to have one of the following (F=fail, N=not Fail) FFFFN FFFNF FFNFF FNFFF NFFFF each has prob \(.3\times.3\times.3\times.3\times.7\) adding them all up gives \(5\times.3\times.3\times.3\times.3\times.7\)